### v1

# New perspectives on the wave equation for a massive scalar field in Schwarzschild spacetime

Abstract
For a massive scalar field in a fixed Schwarzschild background, the radial wave equation obeyed by Fourier modes is first studied. In addition to exact solutions at the regular singular points, we find an exact local solution in the neighbourhood of the irregular singular point at infinity. Moreover, we obtain and study the associated integral representation of the massive scalar field. Eventually, the technique developed for the irregular singular point is applied to the homogeneous equation associated to the inhomogeneous Zerilli equation for gravitational perturbations in a Schwarzschild background.
I Introduction
Almost sixty years after Einstein’s seminal papers Einstein1 ; Einstein2 on general relativity, the study of even just the scalar wave equation in a fixed curved background which solves the Einstein equations was still in the early stages. In the seventies, for example, Persides published very detailed papers on the radial wave equation in Schwarzschild space-time, but claimed that this is not related to any known differential equation of mathematical physics Persides1 ; Persides2 ; Persides3 . Forty years after Persides’ work, the remarkable Ph.D. thesis in Ref. Wardell was still claiming that no exact solution of the radial wave equation is known in Schwarzschild space-time.
It has been therefore our aim to exploit the well-established tools of ordinary and partial differential equations to achieve both exact solution of the radial wave equation for Fourier modes of a massive scalar field, and the integral representation of the full scalar field in Schwarzschild geometry. The scalar field case is already of rich physical interest, in light of modern investigations of null infinity Schmidt ; Kehrberger and of its role in studying Hawking radiation Hawking for quantum fields in curved space-time Figari ; Witten .
The paper is organized as follows. Section II is devoted to an original synthesis of the radial wave equation and its reduction to normal form. Section III obtains approximate solutions of such an equation in the neighbourhood of the two regular singular points and of the irregular singular point. Section IV displays the relation between radial wave equation and the Heun equation. Local solutions at the regular singular points are obtained in Sec. V. A full local solution of the radial wave equation at infinity is obtained in Sec. VI. The resulting integral representation of the massive scalar field that solves the wave equation is studied analytically and numerically in Sec. VII. Section VIII is devoted to the application of Sec. VI to the investigation of the homogeneous equation associated to the inhomogeneous Zerilli equation for gravitational perturbations of a Schwarzschild background. Concluding remarks are presented in Sec. IX, and relevant details are provided in the Appendices.
II Radial wave equation and its reduction to normal form
In agreement with our aims, our starting point is the scalar wave equation
$(\Box-\mu^{2})\psi=0$ (1)
for a scalar field $\psi$ of mass $\mu$ in a fixed Schwarzschild background. This background spacetime is the unique spherically symmetric solution of the vacuum Einstein equations in a way made precise by the Birkhoff theorem HE . In standard spherical coordinates $(t,r,\theta,\phi)$ with $G=c=1$ units, on defining $f(r)=1-\frac{2M}{r}$, the space-time metric is diagonal with components
$g_{\mu\nu}={\rm diag}\Bigr{(}-f,\frac{1}{f},r^{2},r^{2}\sin^{2}\theta\Bigr{)}.$ (2)
Since such components are $t$-independent, one can exploit a Fourier representation according to the expansion
$\psi(t,r,\theta,\phi;\mu)=\sum_{l=0}^{\infty}\psi_{l\mu}(t,r)\sum_{m=-l}^{l}C_{lm}Y_{lm}(\theta,\phi),$ (3)
where $\psi_{l\mu}$ are the integrated Fourier modes
$\psi_{l\mu}(t,r)=\int_{-\infty}^{\infty}e^{-i\omega t}R_{l\omega\mu}(r)\frac{d\omega}{\sqrt{2\pi}},$ (4)
and $R_{l\omega\mu}$ solves the radial wave equation
$\left[\frac{d^{2}}{dr^{2}}+\frac{2(r-M)}{r^{2}f}\frac{d}{dr}+\frac{1}{f}\left(\frac{\omega^{2}}{f}-\frac{l(l+1)}{r^{2}}-\mu^{2}\right)\right]R_{l\omega\mu}(r)=0.$ (5)
This equation can be written in normal form by assuming the product
$R_{l\omega\mu}(r)=\alpha(r)\beta_{l\omega\mu}(r),$ (6)
and looking for an $\alpha(r)$ such that the equation obeyed by $\beta_{l\omega\mu}(r)$ has vanishing coefficient of the first derivative. Indeed, upon setting in Eq. (5)
$p(r)=\frac{2(r-M)}{r^{2}f},\;q_{l\omega\mu}(r)=\frac{1}{f}\left(\frac{\omega^{2}}{f}-\frac{l(l+1)}{r^{2}}-\mu^{2}\right),$ (7)
the general method for achieving normal form Esposito yields
[…]
jointly with the second-order differential equation
$\left[\frac{d^{2}}{dr^{2}}+J_{l\omega\mu}(r)\right]\beta_{l\omega\mu}(r)=0,$ (9)
where
$J_{l\omega\mu}(r)=-\frac{1}{2}\frac{dp}{dr}-\frac{1}{4}p^{2}(r)+q_{l\omega\mu}(r)$ (10)
$=$ $\frac{1}{4r^{2}}+\frac{1}{4(r-2M)^{2}}-\frac{\left(\frac{1}{2}+l(l+1)\right)}{r(r-2M)}$
$+$ $\frac{\omega^{2}r^{2}}{(r-2M)^{2}}-\frac{\mu^{2}r}{(r-2M)}.$
We note that $J_{l\omega\mu}(r)$ has a second-order pole at $r=0$ and at $r=2M$, and hence these are regular singular points, whereas the point $r=\infty$ is an irregular singular point Forsyth as we are going to prove. In order to study the nature of the $r=\infty$ point, one defines the new independent variable $\rho=\frac{1}{r}$, which implies for $\beta_{l\omega\mu}(\rho)$ the differential equation
$\left[\frac{d^{2}}{d\rho^{2}}+\frac{2}{\rho}+\frac{J_{l\omega\mu}(\rho)}{\rho^{4}}\right]\beta_{l\omega\mu}(\rho)=0,$ (11)
where
$\frac{J_{l\omega\mu}(\rho)}{\rho^{4}}=\gamma_{1l\omega}(\rho)+\gamma_{2\omega\mu}(\rho),$ (12)
having defined
$\gamma_{1l\omega}(\rho)=\frac{1}{4\rho^{2}}\left(1+\frac{1}{(1-2M\omega)^{2}}\right)-\frac{\left(\frac{1}{2}+l(l+1)\right)}{\rho^{2}(1-2M\rho)},$ (13)
$\gamma_{2\omega\mu}(\rho)=\frac{1}{\rho^{4}}\left(\frac{\omega^{2}}{(1-2M\rho)^{2}}-\frac{\mu^{2}}{(1-2M\rho)}\right).$ (14)
Since $\gamma_{2\omega\mu}(\rho)$ has a fourth-order pole at $\rho=0$, the point $r=\infty$ is indeed an irregular singular point for Eq. (9).
The material in this section is standard, but nevertheless useful to achieve a self-contained presentation of our subsequent analysis.
III Approximate solutions in the neighbourhood of $r=0,2M$ and $\infty$
III.1 Fourier modes as $r\rightarrow 0$
As $r\rightarrow 0$, the form of $J_{l\omega\mu}(r)$ in Eq. (10) implies that Eq. (9) reduces to a form independent of $l$ and $\omega$, i.e.,
$\left[\frac{d^{2}}{dr^{2}}+\frac{1}{4r^{2}}\right]\beta_{0}(r)=0.$ (15)
On looking for solutions in the form $\beta(r)=r^{\alpha}$, one finds that the square of $\left(\alpha-\frac{1}{2}\right)$ should vanish, and hence there exist two linearly independent integrals giving rise to
$\beta_{0}(r)=C_{1}\sqrt{r}+C_{2}\sqrt{r}\log(r).$ (16)
One requiring regularity at the origin one finds $C_{2}=0$, and hence
$R_{0}(r)=\frac{\beta_{0}(r)}{\sqrt{r(r-2M)}}=\frac{C_{1}}{\sqrt{r-2M}},$ (17)
i.e., the Fourier modes are finite and constant at $r=0$.
III.2 Fourier modes as $r\rightarrow 2M$
As $r\rightarrow 2M$, by virtue of Eq. (10) we find that
$\lim_{r\to 2M}J_{l\omega\mu}(r)=\lim_{r\to 2M}\left[\frac{\chi_{2}(\omega)}{(r-2M)^{2}}+\frac{\chi_{1}(l,\mu)}{(r-2M)}+\chi_{0}\right],$ (18)
where
$\chi_{2}(\omega)=4M^{2}\omega^{2}+\frac{1}{4},$ (19)
$\chi_{1}(l,\mu)=-\frac{1}{2M}\left(\frac{1}{2}+l(l+1)\right)-2M\mu^{2},$ (20)
$\chi_{0}=\frac{1}{16M^{2}}.$ (21)
This implies that Eq. (9) reduces to
$\left[\frac{d^{2}}{dr^{2}}+\frac{\chi_{2}(\omega)}{(r-2M)^{2}}\right]\beta_{\omega,2M}(r)=0,$ (22)
which (up to a multiplicative constant) is solved by
$\beta_{\omega,2M}(r)=(r-2M)^{A},$ (23)
where $A$ is a root of the algebraic equation
$A^{2}-A-\chi_{2}(\omega)=0.$ (24)
Such roots are
$A_{+}=\frac{1}{2}+2iM\omega,$ (25)
$A_{-}=\frac{1}{2}-2iM\omega,$ (26)
and hence the Fourier modes have the limiting form
$R_{\omega,2M}(r)=\frac{\beta_{\omega,2M}(r)}{\sqrt{r(r-2M)}}$ (27)
$=$ $\frac{1}{\sqrt{r}}\Bigr{[}D_{1}(r-2M)^{2iM\omega}+D_{2}(r-2M)^{-2iM\omega}\Bigr{]},$
where $D_{1}$ and $D_{2}$ are suitable constants.
III.3 Fourier modes at infinity
Last but not least, the potential term $J_{l\omega}$ in Eq. (9) has the following Poincaré asymptotic expansion as $r\rightarrow\infty$:
$J_{l\omega\mu}\sim\omega^{2}-\mu^{2}+\frac{A_{1}(\omega,\mu)}{r}+\frac{A_{2}(l,\omega,\mu)}{r^{2}}+\sum_{k=3}^{\infty}\frac{A_{k}(l,\omega,\mu)}{r^{k}}$ (28)
where we have defined
$A_{1}(\omega,\mu)=2M(2\omega^{2}-\mu^{2}),$ (29)
$A_{2}(l,\omega,\mu)=4M^{2}(3\omega^{2}-\mu^{2})-l(l+1).$ (30)
Thus, as $r\rightarrow\infty$, $\beta_{l\omega\mu}(r)$ is (up to a multiplicative constant) the product of the leading order $\beta_{\omega,\infty}(r)=e^{\pm r\sqrt{\mu^{2}-\omega^{2}}}$ times a function that we do not strictly need here, but is part of an analysis that we will perform in Sec. VI.
Note also that, by virtue of Eq. (4), one has to integrate over all values of $\omega$ in order to evaluate the large-$r$ scalar field. Thus, the Fourier modes at large $r$ have to be considered in the following intervals for the $\omega$ variable:
$[-\infty,-\mu],\;]-\mu,\mu[\;{\rm and}\;[\mu,\infty].$
This remark will be exploited in Sec. VII.
IV Connecting the radial with the Heun equation
The reduction to canonical form of Eq. (5) helped us in searching for the singular points $r_{0}=0$, $r_{2M}=2M$ and for the one at infinity. Now we are going to see that Eq. (5) belongs to the class of confluent Heun equations Ronveaux . In order to prove it, we rely upon the change of independent variable frequently used in the literature, i.e.
$z=\frac{(r-r_{2M})}{(r_{0}-r_{2M})}=1-\frac{r}{2M}.$ (31)
In this way, the two finite singularities $r=0,2M$ correspond to $z=1,0$, respectively, and the physical requirement $r\geq 0$ leads to the restriction $z\leq 1$. The equation (5) henceforth takes the form
$\left[\frac{d^{2}}{dz^{2}}-\frac{(1-2z)}{z(1-z)}\frac{d}{dz}+\frac{4M^{2}\omega^{2}(1-z)^{2}}{z^{2}}+\frac{l(l+1)}{z(1-z)}+\frac{4M^{2}\mu^{2}(1-z)}{z}\right]R_{l\omega\mu}(z)=0.$ (32)
Thanks to this change of variable, such equation is nothing but a confluent Heun equation in the so-called non-symmetric canonical form (see Appendix A for a general overview about confluent Heun equations). The non-symmetric canonical form is obtained by inserting in the confluent Heun equation
$H_{{\rm NSCF}}(z)=e^{\alpha_{1}z}\;z^{\alpha_{2}}\;(z-1)^{\alpha_{3}}\;H(z),$ (33)
where $H(z)$ is the confluent Heun function.
V Local solutions at the regular singular points
In our specific case, upon defining the five Heun parameters (no confusion should arise with $\alpha$ and $\beta$ of Sec. II)
$\alpha=4M\sqrt{\mu^{2}-\omega^{2}},$ (34)
$\beta=4iM\omega,$
$\gamma=0,$
$\delta=4M^{2}(\mu^{2}-2\omega^{2}),$
$\eta=-4M^{2}(\mu^{2}-2\omega^{2})-l(l+1),$
Eq. (33) takes the form
$H_{{\rm NSCF}}(z)=e^{\frac{\alpha}{2}z}\;z^{\frac{\beta}{2}}H(z),$ (35)
since $\alpha_{3}=0$ for Eq. (32). Starting from Eq. (35), one can construct the complete Frobenius solution at $z=0$ (corresponding to $r=2M$), that reads (cf. Refs. Fiziev ; Philipp ; Vieira ; Chen ; Minucci )
$R_{0}(z)=e^{\frac{\alpha}{2}z}\;z^{\frac{\beta}{2}}\Bigr{[}C_{1}{\rm HeunC}(\alpha,\beta,\gamma,\delta,\eta,z)+C_{2}z^{-\beta}{\rm HeunC}(\alpha,-\beta,\gamma,\delta,\eta,z)\Bigr{]},$ (36)
whose leading order is
$R_{0}(z)\sim e^{\frac{\alpha}{2}z}\;z^{\frac{\beta}{2}}\Bigr{(}C_{1}+C_{2}z^{-\beta}\Bigr{)}.$ (37)
The Frobenius solution at $z=1$ (i.e., at $r=0$) is given by
$R_{1}(z)$ $=$ $e^{\frac{\alpha}{2}z}\;z^{\frac{\beta}{2}}\Bigr{[}C_{1}{\rm HeunC}(-\alpha,\gamma,\beta,-\delta,\eta+\delta,-z)$ (38)
$+$ $C_{2}(z-1)^{-\gamma}{\rm HeunC}(-\alpha,-\gamma,\beta,-\delta,\eta+\delta,z)\Bigr{]},$
whose leading order is
$R_{1}(z)\sim e^{\frac{\alpha}{2}z}\;z^{\frac{\beta}{2}}\Bigr{(}C_{1}+C_{2}(z-1)^{-\gamma}\Bigr{)}.$ (39)
In order to visualize the results obtained so far, we plot the real (Fig. 1, Fig. 3) and imaginary (Fig. 2, Fig. 4) part of Eqs. (36), (38), with the following choice for the free parameters:
$M=10,\quad\mu=0.1,\quad\omega=1,\quad l=0,\quad C_{1}=C_{2}=1.$ (40)
Figure 1: In this plot, the real part of Eq. (36) is shown for $z\in[-0.01,0.01]$. The $s$-wave near $z=0$ (i.e. $r=2M$) is very close to zero inside the black hole and diverges on the event horizon, then it exhibits strong oscillations.
Figure 2: The imaginary part of Eq. (36) is shown for $z\in[-0.01,0.01]$ The phase of the $s$-wave near $z=0$ behaves qualitatively in the same manner of the real part.
Figure 3: The real part of Eq. (38) is shown for $z\in[0.9,1]$. Near $z=1$ (i.e. $r=0$) the $s$-wave exhibits a plateau, then an overall downward trend as $z$ decreases.
Figure 4: The imaginary part of Eq. (38) is shown for $z\in[0.9,1]$. The phase shows qualitatively the same behaviour of the real part but with sligthly higher values.
VI Full local solution for Fourier modes in the neighbourhood of $r=\infty$
When $r\rightarrow\infty$ one can evaluate $\beta_{l\omega\mu}(r)$, and hence the Fourier modes $R_{l\omega\mu}(r)$, by completing the analysis initiated in Eqs. (28)-(30). For this purpose, we recognize from such equations that the differential equation (9) for $\beta_{l\omega\mu}(r)$ can be written in the form
$\left(\frac{d^{2}}{dr^{2}}+\sum_{k=0}^{\infty}\frac{A_{k}(l,\omega,\mu)}{r^{k}}\right)\beta_{l\omega\mu}(r)=0,$ (41)
where $A_{0}=\omega^{2}-\mu^{2}$, while $A_{1}$ and $A_{2}$ have been already defined in Eqs. (29) and (30), respectively, and all subsequent coefficients can be computed from the asymptotic expansion of Eq. (10) at large $r$. At this stage, Eq. (41) suggests looking for $\beta_{l\omega\mu}(r)$ at large $r$ in the form (setting $\varepsilon=\pm 1$ and considering $\zeta$ as a parameter to be determined as shown below)
$\beta_{l\omega\mu}(r)\sim e^{\varepsilon r\sqrt{-A_{0}}}\;r^{\zeta}\;\left(1+\sum_{s=1}^{\infty}\frac{B_{s}(l,\omega,\mu)}{r^{s}}\right).$ (42)
The insertion of Eq. (42) into Eq. (41) yields
$\gamma_{0}+\frac{\gamma_{1}}{r}+\frac{\gamma_{2}}{r^{2}}+\sum_{k=3}^{\infty}\frac{\gamma_{k}}{r^{k}}=0.$ (43)
For this equation to be identically satisfied as $r\rightarrow\infty$, all $\gamma_{k}$ coefficients should vanish. Indeed, we find
$\gamma_{0}=A_{0}-A_{0}=0,$ (44)
while
$\gamma_{1}=2\zeta\varepsilon\sqrt{-A_{0}}+A_{1},$ (45)
$\gamma_{2}=(\zeta-1)(2\varepsilon B_{1}\sqrt{-A_{0}}+\zeta)+A_{1}B_{1}+A_{2},$ (46)
jointly with infinitely many other equations for all subsequent $\gamma_{k}$ coefficients. Since they should all vanish, we obtain linear algebraic equations for $B_{1},B_{2},...,B_{\infty}$. In addition, since we look for a bounded solution at large $r$, we restrict ourselves to the $\varepsilon=-1$ case. For instance, from Eq. (45) we evaluate the explicit form of $\zeta$, while from Eq. (46) we obtain $B_{1}$, i.e.
$\zeta=\frac{A_{1}}{2\sqrt{-A_{0}}}=\frac{M(2\omega^{2}-\mu^{2})}{\sqrt{\mu^{2}-\omega^{2}}},$ (47)
$B_{1}=\frac{\zeta(\zeta-1)+A_{2}}{2(\zeta-1)\sqrt{-A_{0}}-A_{1}}.$ (48)
By virtue of Eq. (6), we find eventually the large-$r$ Fourier modes in the form (cf. Refs. Fiziev ; Philipp ; Minucci )
$R_{l\omega\mu,\infty}(r)\sim\frac{e^{-r\sqrt{-A_{0}}}}{\sqrt{r(r-2M)}}r^{\frac{A_{1}}{2\sqrt{-A_{0}}}}\left(1+\sum_{s=1}^{\infty}\frac{B_{s}(l,\omega,\mu)}{r^{s}}\right).$ (49)
Note that, by considering only the $k=0,1,2$ terms in Eq. (41), one would obtain the Whittaker equation (a modified form of the confluent hypergeometric equation), solved by the two Whittaker functions. However, all negative powers of $r$ occur in Eq. (41), and hence one needs our factorization in Eq. (49), where the exponential cancels the effect of $A_{0}$, and the $B_{s}$ functions of $l,\omega$ compensate the negative powers of $r$ in the operator provided that a non-integer power of $r$ is included in our factorized ansatz, in agreement with the general method of Poincaré for linear differential equations Poincare ; Poincare2 .
In Appendix C we offer a theoretical motivation for the ansatz given by Eq. (42) and the explicit computation of $\zeta$ and the first $A_{k}$, $B_{s}$ coefficients. In order to visualize such a solution, we consider Eq. (49) with the same choice for parameters given by Eq. (40) (this time we don’t have integration constants like $C_{1}$ and $C_{2}$), ending up with two plots: one for the real part (Fig. 5) of the field and the other for the imaginary one (Fig. 6). The relation with the local results by Persides in the massless case Persides1 is discussed in Appendix D.
Figure 5: The real part of Eq. (49) is plotted with parameters given by Eq. (40) summing over $s$ from $1$ to $5$ for $r\in[10M,15M]$. The oscillations of the field become smaller and smaller as $r$ grows.
Figure 6: The imaginary part of Eq. (49) is plotted with parameters given by Eq. (40) summing over $s$ from $1$ to $5$ for $r\in[10M,15M]$. The phase follows qualitatively the same behaviour of the real part.
VII Large-$r$ solution of the full wave equation
So far, we have studied the behaviour of solutions for fixed frequencies. Since one of our goals is to study the spectrum of a massive field far away from the black hole, we now focus the attention on the study of the solution at spacelike infinity by including the contribution of all frequencies (neglecting the angular part). In the previous sections we have discovered that, when $r\rightarrow\infty$, the Fourier modes take the approximate form (49). The large-$r$ form of the scalar field $\psi$ which solves Eq. (1) and satisfies the asymptotic condition
$\lim_{r\to\infty}|\psi_{l}(t,r)|<\infty$ (50)
is therefore expressed by Eq. (4), where, upon defining
$\varphi(t,r,\omega)=\omega t+r\sqrt{\omega^{2}-\mu^{2}}+M\frac{(2\omega^{2}-\mu^{2})}{\sqrt{\omega^{2}-\mu^{2}}}\log(r),$ (51)
we find, for $\omega^{2}>\mu^{2}$,
$\psi^{I}_{l\mu}(t,r)\sim\int_{-\infty}^{\infty}e^{-i\omega t}R_{l\omega\mu,\infty}(r)\frac{d\omega}{\sqrt{2\pi}}$ (52)
$\sim$ $\frac{1}{\sqrt{r(r-2M)}}\left(\int_{-\infty}^{-\mu}+\int_{\mu}^{\infty}\right)\frac{d\omega}{\sqrt{2\pi}}e^{-i\varphi(t,r,\omega)}\left(1+\sum_{s=1}^{\infty}\frac{B_{s}(l,\omega,\mu)}{r^{s}}\right).$
The integral is qualitatively different in the $\omega^{2}<\mu^{2}$ case, since the last two terms in Eq. (51) become imaginary, leading to
$\psi^{II}_{l\mu}(t,r)\sim\frac{1}{\sqrt{r(r-2M)}}\int_{-\mu}^{\mu}\frac{d\omega}{2\pi}e^{-i\omega t}e^{r\sqrt{\mu^{2}-\omega^{2}}\left(1+M\frac{\mu^{2}-2\omega^{2}}{\sqrt{\mu^{2}-\omega^{2}}}\frac{\log(r)}{r}\right)}\left(1+\sum_{s=1}^{\infty}\frac{B_{s}(l,\omega,\mu)}{r^{s}}\right).$ (53)
The investigation of these two integrals is a hard task. However, we can rely on analytical approximations and numerical analysis in order to extract the essential features from them. The starting point is given by studying the behaviour in frequencies of the two integrand functions, which we label as
$\Omega_{l\mu}^{I}(t,r,\omega)=\frac{1}{\sqrt{2\pi r(r-2M)}}e^{-i\varphi(t,r,\omega)}\left(1+\sum_{s=1}^{\infty}\frac{B_{s}(l,\omega\mu)}{r^{s}}\right),$ (54)
$\Omega_{l\mu}^{II}(t,r,\omega)=\frac{1}{\sqrt{2\pi r(r-2M)}}e^{-i\omega t}e^{r\sqrt{\mu^{2}-\omega^{2}}\left(1+M\frac{\mu^{2}-2\omega^{2}}{\sqrt{\mu^{2}-\omega^{2}}}\frac{\log(r)}{r}\right)}\left(1+\sum_{s=1}^{\infty}\frac{B_{s}(l,\omega,\mu)}{r^{s}}\right).$ (55)
Notice that the $l$-dependence is encoded in the coefficients of the series expansion, requiring to add also the angular part (given by spherical harmonics) to study the superposition of several waves. We therefore limit ourselves to the study of the $s$-wave, given by the constraint $B_{s}(l,\omega,\mu)=B_{s}(0,\omega,\mu)$. Moreover, for numerical reasons, we consider only the first five negative powers of $r$, arriving at a semplified version of Eqs. (54), (55). Upon setting
$\mu=0.1,\quad t=1,\quad M=10,$ (56)
we first consider the plots of the real (Fig. 7) and imaginary (Fig. 8) part of Eq. (55).
Figure 7: Plot of Re$\left[\Omega_{0\mu}^{II}(1,r,\omega)\right]$ for $r\in[100,120]$, $\omega\in[-\mu,\mu]$, setting $\mu=0.1$ and $M=10$, up to the $r^{-5}$ term. Along the $\omega$-axis the function has a Gaussian-like behaviour (it is possible to prove it by expanding in series Eq. 55 for small frequencies), while it diverges along the $r$-axis. The integral, which is of course the area under the curve, is clearly divergent in the $r\rightarrow\infty$ limit.
Figure 8: Plot of Im$\left[\Omega_{0\mu}^{II}(1,r,\omega)\right]$ for $r\in[100,120]$, $\omega\in[-\mu,\mu]$, setting $\mu=0.1$ and $M=10$, up to the $r^{-5}$ term. Unlike the real part, the phase shows a different behaviour by virtue of its odd-parity with respect to $\omega$. The contributions to the integral on the $r$-axis seem to vanish in light of this feature.
It is possible to evaluate numerically the integral given by Eq. (53) by expanding $\Omega_{l\mu}^{II}(t,r,\omega)$ around $\omega=0$, since this region gives the largest contribution to it (as shown in Fig. 7). We have performed such evaluation up to the sixth power of $\omega$ for the choosen parameters, arriving at the conclusion that such a piece diverges. In order to obtain also an analytical estimate of such a divergence, we consider the change of variable given by
$\omega=\mu\sin(\tau)$ (57)
applied to the leading term of Eq. (55), obtaining therefore
$\Omega_{l\mu}^{II}(t,r,\tau)\approx\frac{1}{\sqrt{2\pi r(r-2M)}}r^{\mu M(2\left|\cos(\tau)\right|-\text{sign}(\cos(\tau))\sec(\tau))}e^{\mu(r\left|\cos(\tau)\right|-i\sin(\tau))}.$ (58)
The requirement $\omega\ll 1$ yields $\tau\ll 1$, giving as leading term
$\Omega_{l\mu}^{II}(t,r,\tau)\approx\frac{1}{\sqrt{2\pi r(r-2M)}}r^{\mu M}e^{\mu r}(i\mu t\tau-1)+\text{O}(\tau^{2}).$ (59)
After a straightforward integration we arrive at the desired result, i.e.,
[…]
Even though it is possible to perform more accurate computations (including more terms and/or the negative powers of $r$ of the series expansion) it is clear that the dominant contribution is the one of the exponential term, which causes the divergence of the whole integral in the $r\rightarrow\infty$ limit, as shown by the purely numerical analysis given before. Notice that the field mass term, $\mu$, plays a central role: in the massless case the equivalent of such a piece vanishes (as it happens here in the $\mu\rightarrow 0$ limit). We have therefore to face a case where a field not only does not vanish at spacelike infinity, but furthermore it diverges exponentially. As far as we can see, this property is compatible with the findings in Ref. Huh , where part (II) of Proposition 2.1 therein shows that the norm of a massive scalar field obeying the wave equation in Schwarzschild spacetime is majorized by the product of two functions, one of which is indeed divergent in a suitable limit. This agreement merits explicit mention, because the work in Ref. Huh has exploited advanced methods from the modern theory of hyperbolic equations, whereas we have used the standard techniques of classical mathematical physics.
The case of $\Omega_{l\mu}^{I}(t,r,\omega)$ is strongly different. Upon relying on the same conventions, we show the plots of its real part (the ones for the phase are very similar) for $\omega\in[-10^{3},-\mu]$ in Fig. 9 and for $\omega\in[\mu,10^{3}]$ in Fig. 10.
Figure 9: Plot of Re$\left[\Omega_{0\mu}^{I}(1,r,\omega)\right]$ for $r\in[1,1.2]\cdot 10^{3}$, $\omega\in[-10^{3},-\mu]$, setting $\mu=0.1$ and $M=10$, up to the $r^{-5}$ term. For every fixed value of $r$, the field exhibits increasing oscillations as $\omega$ grows (in this specific case for $\omega<-400$. In the low-frequency limit, the field is neglibile. Both behaviours suggest that the numerical interval is not divergent.
Figure 10: Plot of Re$\left[\Omega_{0\mu}^{I}(1,r,\omega)\right]$ for $r\in[1,1.2]\cdot 10^{3}$, $\omega\in[\mu,10^{3}]$, setting $\mu=0.1$ and $M=10$, up to the $r^{-5}$ term. The discussion of such a figure follows closely the one of Fig. 9.
On both numerical and analytical ground we can say that, while the function $\Omega_{l}^{I}(t,r,\omega)$ diverges, the associated integral (given by Eq. (52)) remains finite and eventually vanishes at spacelike infinity. In fact, we have an oscillating term $e^{-i\varphi}$ alleviated by an infinite series of negative powers of $r$.
Interestingly, we find that a massive scalar field contains a term which diverges at spacelike infinity, while this feature is absent in the massless case. In conclusion, the mass plays a crucial role in the asymptotic description of a scalar field in spherical coordinates.
VIII Application of our Sec. VI to the Zerilli equation
The systematic algorithm of Sec. VI can be applied to equations of even greater interest. For this purpose, we here consider the recent analysis in Ref. Hamaide of the Zerilli equation Zerilli , which is obtained by perturbing a background Schwarzschild metric with a small parity-even perturbation expressed by a symmetric tensor field of type $(0,2)$. Eventually, the work in Ref. Hamaide obtains an inhomogeneous radial wave equation, that we express in the form
$\left[\frac{d^{2}}{dr^{2}}+p(r)\frac{d}{dr}+q(r,l,\omega)\right]u(r,l,\omega)={\widetilde{I}}(r,l,\omega),$ (61)
where
$p(r)=\frac{2M}{r(r-2M)},$ (62)
$q(r,l,\omega)=\frac{r^{2}}{(r-2M)^{2}}(\omega^{2}-V(r,l,\omega)),$ (63)
having defined
$V(r,l,\omega)=\left(1-\frac{2M}{r}\right)\frac{[2\eta^{2}(\eta+1)r^{3}+6\eta^{2}Mr^{2}+18\eta M^{2}r+18M^{3}]}{r^{3}(\eta r+3M)^{2}},$ (64)
where $\eta=\frac{(l-1)(l+2)}{2}$. The work in Ref. Hamaide uses a Green-function method in order to solve the inhomogeneous differential equation (61) by using the well-established technique for obtaining the Green function of the operator on the left-hand side of Eq. (61) from solutions of the associated homogeneus equation
$\left[\frac{d^{2}}{dr^{2}}+p(r)\frac{d}{dr}+q(r,l,\omega)\right]U(r,l,\omega)=0.$ (65)
It has been our choice to focus on this particular aspect of the general problem, since we have already written several sections on the massive scalar case, which is the main topic of our paper.
By repeating the reduction to normal form as we did in Sec. II, we obtain $U(r,l,\omega)$ in the factorized form
$U(r,l,\omega)=\gamma(r)\beta(r,l,\omega),$ (66)
where
$\gamma(r)=\frac{1}{\sqrt{1-\frac{2M}{r}}},$ (67)
while $\beta(r,l,\omega)$ solves the equation
$\left[\frac{d^{2}}{dr^{2}}+J(r,l,\omega)\right]\beta(r,l,\omega)=0,$ (68)
having defined
$J(r,l,\omega)=\frac{1}{(r-2M)^{2}r^{2}[(l(l+1)-2)r+6M]^{2}}\Bigr{[}\omega^{2}(l-1)^{2}(l+2)^{2}r^{6}$ (69)
$+$ $12M\omega^{2}(l-1)(l+2)r^{5}+(-l^{6}-3l^{5}+l^{4}+7l^{3}-4l+36M^{2}\omega^{2})r^{4}$
$+$ $2M(l-1)^{3}(l+2)^{3}r^{3}+9M^{2}\left(l(l+1)-\frac{10}{3}\right)(l+2)(l-1)r^{2}$
$+$ $36M^{3}(l-1)(l+2)r+36M^{4}\Bigr{]}.$
At this stage, in close analogy with Eq. (41), we may expand Eq. (68) as $r\rightarrow\infty$ in the form
$\left(\frac{d^{2}}{dr^{2}}+\sum_{k=0}^{\infty}\frac{A_{k}(l,\omega)}{r^{k}}\right)\beta(r,l,\omega)=0,$ (70)
where the function $\beta$ has the asymptotic expansion
$\beta(r,l,\omega)\sim e^{-r\sqrt{-A_{0}}}\;r^{\zeta}\left(1+\sum_{s=1}^{\infty}\frac{B_{s}(l,\omega)}{r^{s}}\right).$ (71)
The formulae for expressing $\zeta$ and $B_{s}$ in terms of the $A_{k}(l,\omega)$ coincide with the ones in Sec. VI, whereas the $A_{k}$ are different. Indeed, we find for example that
$A_{0}=\omega^{2},$ (72)
$A_{1}=4\omega^{2}M,$ (73)
$A_{2}=12M^{2}\omega^{2}-l(l+1),$ (74)
$A_{3}=\frac{2M[16M^{2}\omega^{2}(l(l+1)-2)-l^{4}-2l^{3}+5l^{2}+6l+4]}{l(l+1)-2},$ (75)
and hence, from Eqs. (47) and (48),
$\zeta=-2iM|\omega|,$ (76)
$B_{1}=-M+\frac{i[8M^{2}\omega^{2}-l(l+1)]}{2|\omega|},$ (77)
supplemented by a countable infinity of higher-order $A_{k}$ and $B_{s}$ coefficients.
Hereafter, we plot the real (Fig. 11) and imaginary (Fig. 12) part of Eq. (66) upon considering the parameters
$M=10,\quad\omega=1,\quad l=0,$ (78)
up to the $B_{5}$ coefficient in Eq. (71). Our findings in Figs. 11 and 12 are in full agreement with the plane-wave behaviour as $r\rightarrow\infty$ obtained in Ref. Hamaide .
Figure 11: The real part of Eq. (66) is shown for $r\in[100,300]$, given the parameters (78) and summing over $s$ from 1 to 5 in Eq. (71). After achieving high values in the region closest to the event horizon, the field shrinks at $r\approx 140$ and then behaves like a plane wave, as it should be.
Figure 12: The imaginary part of Eq. (66) is shown for $r\in[100,300]$, given the parameters (78) and summing over $s$ from 1 to 5 in Eq. (71). Its behaviour is very similar to the one of the real part, leading to the same conclusions.
IX Concluding remarks and open problems
The scalar wave equation in Schwarzschild spacetime has been studied intensively over more than half a century Persides1 ; Persides2 ; Persides3 ; Schmidt ; Fiziev ; Philipp ; Kehrberger , until the very recent and highly valuable work in Ref. Bianchi , but the material in our Secs. VI and VII remains original. More precisely, as far as we know, our original contributions are as follows: (i) For a massive scalar field in Schwarzschild geometry, our Eqs. (47)-(49) and (108)-(109) yield the full Poincaré asymptotic expansion of Fourier modes that solve the radial wave equation. (ii) Our Eqs. (52)-(53) yield the integral representation of the massive scalar field, with the resulting numerical analysis discussed therein.
For massive scalar fields in Schwarzschild spacetime, the literature has mainly focused on late-time behaviour Burko of Fourier modes, whereas rigorous results on solutions of the Cauchy problem are along the lines of Ref. Huh . We think that the next step will consist in applying analogous techniques to other spherically symmetric spacetimes, both in black-hole theory Minucci and in cosmological backgrounds BE . For this purpose, we have eventually studied the homogeneous equation associated to the inhomogeneous Zerilli equation in a Schwarzschild background. Our original formulae (72)-(77) can be applied to build the Green function, which in turn leads to the solution of the inhomogeneous Zerilli equation as shown in Ref. Hamaide .
Acknowledgements
G.E. is grateful to INDAM for membership.
Appendix A The Confluent Heun equation and its solutions
Given a differentiable function $Y$ of the variable $z$, the confluent Heun equation Ronveaux is a second-order ordinary differential equation of the following form:
$\left[\frac{d^{2}}{dz^{2}}+\left(\alpha+\frac{(\beta+1)}{z}+\frac{(\gamma+1)}{(z-1)}\right)\frac{d}{dz}+\left(\frac{\mu}{z}+\frac{\nu}{(z-1)}\right)\right]Y(z)=0,$ (79)
where the parameters $\mu$ and $\nu$ obey the relations
$\mu=\frac{1}{2}(\alpha-\beta-\gamma-2\eta+\alpha\beta-\beta\gamma),$ (80)
$\nu=\frac{1}{2}(\alpha+\beta+\gamma+2\delta+2\eta+\alpha\gamma+\beta\gamma),$ (81)
while $\alpha,\beta,\gamma,\delta,\eta$ are free (in general complex) parameters. A Heun equation has four singular points, while its confluent form has two regular singular points at $z=0$, $z=1$ of $s$-rank $1$ and an irregular singular point at $z=\infty$ of $s$-rank $2$. Indeed, the confluent form is obtained from the standard one by glueing two singularities. In the literature, only power-series solutions in the neighbourhood of singular points are known, resulting in three sets of two linearly independent solutions. The ones in the neighbourhood of $z=0$ and $z=1$ are called Frobenius solutions, while the ones in the neighbourhood of the point at infinity are known as Thomé solutions. Upon denoting with
${\rm HeunC}(\alpha,\beta,\gamma,\delta,\eta,z)$
the first Frobenius solution as $z$ approaches 00, we have that
$Y_{0}(z)=C_{1}{\rm HeunC}(\alpha,\beta,\gamma,\delta,\eta,z)+C_{2}z^{-\beta}{\rm HeunC}(\alpha,-\beta,\gamma,\delta,\eta,z)$ (82)
is the complete Frobenius solution as $z\rightarrow 0$, while
$Y_{1}(z)=C_{1}{\rm HeunC}(-\alpha,\gamma,\beta,-\delta,\eta+\delta,1-z)$ (83)
$+$ $C_{2}(z-1)^{-c}{\rm HeunC}(-\alpha,-\gamma,\beta,-\delta,\eta+\delta,1-z)$
is the complete Frobenius solution as $z\rightarrow 1$. Their leading order is given by
$Y_{0}(z)\sim C_{1}+C_{2}z^{-\beta},$ (84)
$Y_{1}(z)\sim C_{1}+C_{2}(z-1)^{-c}.$ (85)
The function ${\rm HeunC}$ is implemented in many symbolic computer programs such as Mathematica and Maple. In our paper we are adopting Maple conventions. Once that ${\rm HeunC}$ is known, it is straightforward to build the two sets of linearly independent Frobenius solutions written before. Their radius of convergence is obtained from the conditions that the absolute value of the argument of ${\rm HeunC}$ must be less than $1$: $|z|<1$. The solution for all values of $z$ is obtained by overlapping the analytical continuations of $Y_{0}(z)$ and $Y_{1}(z)$.
The solution at infinity is quite different. It is well known that every Heun equation (also the confluent, biconfluent and three-confluent one) admits a solution at infinity whose asymptotic expansion is given by
$Y_{\infty}(z)=e^{T_{n}(z)}\;z^{-\theta}\;\sum_{k=0}^{\infty}\rho_{k}z^{-k},$ (86)
where $T_{n}$ is a polynomial function of $z$. Note that the sign of $\theta$ may change, depending on the chosen parametrization. Upon setting $T_{n}(z)\sim 1+z$ one obtains the so-called Thomé solutions
$Y_{\infty}(z)=C_{1}z^{-\theta_{1}}\sum_{k=0}^{\infty}\rho_{k}z^{-k}+C_{2}e^{\theta_{2}z}\;z^{-\theta_{3}}\;\sum_{k=0}^{\infty}\sigma_{k}z^{-k}.$ (87)
By inserting them into Eq. (79) it is possible to obtain a recurrence relation for $\rho_{k},\sigma_{k}$. Note that $\theta_{1},\theta_{2},\theta_{3}$ are in principle different and complex. However, it is possible to relate them to $\alpha,\beta,\gamma,\delta,\eta$ even though this strongly depends on the chosen parametrization.
Appendix B Canonical reduction with respect to $z$
It is very easy to perform the canonical reduction of the radial equation (31) with respect to the independent variable $z$. Following the same procedure of Sec. II, we obtain
$\alpha(z)=\frac{1}{\sqrt{z(1-z)}},$ (88)
while the potential term reads as
$T_{l\omega\mu}(z)=\frac{(1+16M^{2}\omega^{2})}{4z^{2}(z-1)^{2}}+\frac{(4M^{2}(\mu^{2}-4\omega^{2})+l(l+1))}{z(z-1)^{2}}$ (89)
$+$ $\frac{(12M^{2}(2\omega^{2}-\mu^{2})-l(l+1))}{(z-1)^{2}}+\frac{4M^{2}(3\mu^{2}-4\omega^{2})z}{(z-1)^{2}}$
$+$ $\frac{4M^{2}(\omega^{2}-\mu^{2})z^{2}}{(z-1)^{2}}.$
The function $\beta_{l\omega}$ is the solution of the differential equation
$\left[\frac{d^{2}}{dz^{2}}+T_{l\omega\mu}(z)\right]\beta_{l\omega\mu}(z)=0.$ (90)
In the neighbourhood of $z=0$, the potential takes the approximate form
$T_{0l\omega}(z)\sim\left(4M^{2}\omega^{2}+\frac{1}{4}\right)z^{-2},$ (91)
which leads to the approximate solution
$\beta_{0l\omega}(z)=D_{1}z^{2iM\omega+\frac{1}{2}}+D_{2}z^{-2iM\omega+\frac{1}{2}}.$ (92)
In the neighbourhood of $z=1$ we find instead (after imposing regularity of $\beta$ therein)
$T_{1l\omega}(z)\sim\frac{1}{4}(z-1)^{-2},$ (93)
$\beta_{1l\omega}(z)=C_{1}\sqrt{z-1}.$ (94)
At infinity (i.e., as $r$ approaches +$\infty$+ ∞, which implies that $z$ approaches $\infty$- ∞) we find
$T_{\infty l\omega\mu}(z)\sim 4M^{2}(\omega^{2}-\mu^{2}),$ (95)
$\beta_{\infty l\omega\mu}(z)=S_{1}e^{2zM\sqrt{\omega^{2}-\mu^{2}}}+S_{2}e^{-2z\sqrt{\omega^{2}-\mu^{2}}}.$ (96)
We note complete correspondence with the leading behaviours obtained in Sec. III with respect to the $r$ variable.
Appendix C Remarks on the solution at infinity
First, we provide a theoretical derivation of the ansatz given by Eq. (42) based on a method developed by Poincaré Poincare ; Poincare2 to compute the solution of a differential equation with variable coefficients. We then evaluate the explicit form of the first coefficients $B_{k}$ of the solution given by Eq. (49).
C.1 Theoretical derivation
The radial wave equation in its normal form (9), with the potential term defined in Eq. (10), can be solved locally, i.e., in the neighbourhhod of the singular points Persides1 ; Fiziev ; Philipp ; Minucci . For this purpose, it is helpful to re-express it in a form where the variable coefficients are polynomials, following the seminal work of Poincaré Poincare . This is achieved by multiplying it on the left by $r^{2}(r-2M)^{2}$. This leads to the differential equation
$\left[P_{2}(r)\frac{d^{2}}{dr^{2}}+P_{0l\omega\mu}(r)\right]\beta_{l\omega\mu}(r)=0,$ (97)
where
$P_{2}(r)=r^{2}(r-2M)^{2}=r^{4}-4Mr^{3}+4M^{2}r^{2},$ (98)
$P_{0l\omega\mu}(r)=r^{2}(r-2M)^{2}J_{l\omega\mu}(r)$ (99)
$=$ $(\omega^{2}-\mu^{2})r^{4}+2M\mu^{2}r^{3}+\left[\frac{1}{2}-\frac{1}{2}-l(l+1)\right]r^{2}$
$+$ $\Bigr{[}-M+M+2Ml(l+1)\Bigr{]}r+M^{2}$
$=$ $(\omega^{2}-\mu^{2})r^{4}+2M\mu^{2}r^{3}-l(l+1)r^{2}+2Ml(l+1)r+M^{2},$
and we have stressed in the intermediate step the exact cancellations occurring in the evaluation of the polynomial (99). As far as we know, only the authors of Ref. Philipp came pretty close to studying an equation like our (97) with $r$ as independent variable, but they chose a third degree polynomial in their Eq. (32) as the coefficient of the second derivative of the desired solution.
At this stage, the approximate solutions (17), (27) and the one which would be obtained for Eq. (28) by considering only the leading term (labelled as $\beta_{\omega,\infty}$) suggest looking for a solution of Eq. (97) in the factorized form of product (called $s$-homotopic ansatz in the current literature Philipp ; Minucci )
$\beta_{l\omega\mu}(r)=\beta_{0}(r)\beta_{\omega,2M}(r)\beta_{\omega,\infty}(r)G_{l\omega\mu}(r),$ (100)
where the function $G_{l\omega\mu}$ should be such that this product solves, in the neighbourhood of $r=0,2M,\infty$, Eq. (97). In the ansatz (100) we have
$\beta_{0}(r)\beta_{\omega,2M}(r)\beta_{\omega,\infty}(r)=\sqrt{r(r-2M)}\Bigr{[}D_{1}(r-2M)^{2iM\omega}+D_{2}(r-2M)^{-2iM\omega}\Bigr{]}$ (101)
$\times$ $\Bigr{[}S_{1}e^{r\sqrt{\mu^{2}-\omega^{2}}}+S_{2}e^{-r\sqrt{\mu^{2}-\omega^{2}}}\Bigr{]}.$
The factorization in Eq. (100) follows what was learned from Poincaré Poincare2 , i.e., for linear differential equations with variable coefficients, the solution consists of the limiting form at the origin times the limiting form at infinity times a suitable series. With respect to the case considered in Ref. Poincare2 , we are here dealing with two (rather than just one) regular singular points at finite values of the independent variable. The structure of the ansatz given by Eqs.(100), (101) is too rich and it does not provide any semplification of computations. Our proposal, given by Eq. (42), is an easier case where $S_{1}=0$, $\beta_{0}(r)\beta_{\omega,2M}(r)\simeq r^{\alpha}$ and the interpolating series $F_{l\omega}(r)$ consists only of negative powers of r. In this way we are able to study the behaviour of the field at infinity without any effort.
C.2 Some coefficients
We begin by writing the first coefficients $A_{k}$ of the asymptotic expansion of the canonical potential at infinity, given by Eq. (28)
$A_{0}=\omega^{2}-\mu^{2},$ (102)
$A_{1}=2M(2\omega^{2}-\mu^{2}),$ (103)
$A_{2}=4M^{2}(3\omega^{2}-\mu^{2})-l(l+1),$ (104)
$A_{3}=2M[16\omega^{2}M^{2}-4\mu^{2}M^{2}-l(l+1)],$ (105)
$A_{4}=M^{2}[80\omega^{2}M^{2}-16\mu^{2}M^{2}-4l(l+1)+1],$ (106)
$A_{5}=4M^{3}[48M^{2}\omega^{2}-8\mu^{2}M^{2}-2l(l+1)+1].$ (107)
In ligth of the above, the first two $B_{s}$ coefficients are explicitly obtained here:
$\begin{split}B_{1}=\frac{1}{2(\mu^{2}-\omega^{2})^{2}}&\biggr{\{}\sqrt{\mu^{2}-\omega^{2}}\Bigr{[}3M^{2}\mu^{4}+\mu^{2}(l(l+1)-12\omega^{2}M^{2})+8M^{2}\omega^{4}\\ &-l(l+1)\omega^{2}\Bigr{]}-M\mu^{4}+3M\mu^{2}\omega^{2}-2M\omega^{4}\biggr{\}},\end{split}$ (108)
$\begin{split}B_{2}=&\frac{1}{2(\mu^{2}-\omega^{2})^{3}}\biggr{\{}4M\sqrt{\mu^{2}-\omega^{2}}\Bigr{[}\omega^{4}(4M^{2}\mu^{2}-l(l+1)+1)\\ &-6\omega^{2}\mu^{2}\left(M^{2}\mu^{2}-\frac{1}{6}l(l+1)+\frac{1}{4}\right)+M^{2}\mu^{6}+\frac{\mu^{4}}{2}\Bigr{]}\\ &+64M^{4}\omega^{8}-192M^{2}\omega^{6}\left(M^{2}\mu^{2}+\frac{1}{12}l(l+1)-\frac{1}{48}\right)\\ &+\omega^{4}\Bigr{[}192M^{4}\mu^{4}+40M^{2}\mu^{2}\left(l(l+1)-\frac{2}{5}\right)+l^{4}+2l^{3}-l^{2}-2l\Bigr{]}\\ &-72\mu^{2}\omega^{2}\Bigr{[}M^{4}\mu^{4}+\frac{5}{12}M^{2}\mu^{2}\left(l(l+1)-\frac{1}{2}\right)+\frac{1}{36}l(l+2)(l^{2}-1)\Bigr{]}\\ &+9\mu^{4}\left(M^{2}\mu^{2}+\frac{1}{3}l^{2}-\frac{1}{3}\right)\left(M^{2}\mu^{2}+\frac{l^{2}}{3}+\frac{2}{3}l\right)\biggr{\}}.\end{split}$ (109)
For $s>2$, the coefficients start getting cumbersome. However, their evaluation remains feasible by virtue of the algorithm described in the main text.
Appendix D Comparison with the Persides analysis
Sections VI and VII have presented our original calculations for a massive scalar field in a fixed Schwarzschild background, whereas, for a massless field, the large-$r$ solution of the radial wave equation was first obtained by Persides Persides1 by using a $3$-term recurrence relation, because he did not reduce the equation for Fourier modes $R_{l\omega}$ to its normal form. With our notation, the Persides large-$r$ form of $R_{l\omega}(r)$ reads as (setting $c$ and $G$ to $1$ in Ref. Persides1 )
$R_{l\omega,\pm}(r)\sim e^{\mp i\omega(r+2M\log(r-2M))}F_{l\omega,\pm}(r),$ (110)
the function $F_{l\omega,\pm}$ being obtainable as
$F_{l\omega,\pm}(r)=F_{l}(\omega r;\epsilon=\pm 2i),$ (111)
where $F_{l}(\omega r;\epsilon)$ solves, by definition, the differential equation (here $z=\omega r$)
$\biggr{[}z(z-2M\omega)\frac{d^{2}}{dz^{2}}+(-\epsilon z^{2}+2z-2M\omega)\frac{d}{dz}-(\epsilon z+l(l+1))\biggr{]}F_{l}=0,$ (112)
and has the large-$z$ (i.e., large-$r$) asymptotic expansion
$F_{l}(\omega r;\epsilon)\sim\left(\frac{\epsilon}{2}\right)^{l+1}\sum_{n=0}^{\infty}\tau_{n}(l,\omega)z^{-(n+1)},$ (113)
where $\tau_{0}=1$, while Persides1
$n\epsilon\tau_{n}-(l+n)(l-n+1)\tau_{n-1}-2M\omega(n-1)^{2}\tau_{n-2}=0.$ (114)
The explicit form of the first four $\tau_{n}$ obtained by such a relation in the general $\tau_{0}\neq 1$ case is as follows:
$\begin{split}\tau_{1}=&\frac{\tau_{0}}{\epsilon}(l+1)l,\\ \tau_{2}=&\frac{\tau_{0}}{2\epsilon^{2}}[(l+2)(l^{2}-1)l+2M\omega\epsilon],\\ \tau_{3}=&\frac{\tau_{0}}{6\epsilon^{3}}[(l+3)(l^{5}-5l^{3}+4l)+6M\omega\epsilon(3l^{2}+3l-2)],\\ \tau_{4}=&\frac{\tau_{0}}{24\epsilon^{4}}[(l+4)(l^{7}-14l^{5}+49l^{3}-36l)+12M\omega\epsilon(6l^{4}+9M\omega\epsilon+12l^{3}-22l^{2}-28l+12)].\end{split}$
On the other hand, in light of our Sec. VI, one might be tempted to study the massless case by defining
$R_{l\omega,+}(r)\sim\frac{1}{\sqrt{r(r-2M)}}e^{-i\omega(r+2M\log(r))}\left(1+\sum_{s=1}^{\infty}\frac{B_{s}(l,\omega,0)}{r^{s}}\right).$
However, such a formula cannot be taken to hold, because Eq. (49) is only a Poincaré asymptotic expansion (see Appendix E), and hence
$\lim_{\mu\to 0}\sum_{s=1}^{\infty}\frac{B_{s}(l,\omega,\mu)}{r^{s}}\not=\sum_{s=1}^{\infty}\lim_{\mu\to 0}\frac{B_{s}(l,\omega,\mu)}{r^{s}}.$
As far as we can see, the best that one can do is to check that, by including as many $B_{s}(l,\omega,\mu)$ coefficients as possible in the evaluation of the massive case, and as many $\tau_{n}(l,\omega)$ coefficients as possible in the massless case, one obtains an accurate approximation of the local solution of the radial wave equation in the neighbourhood of $r=\infty$. In the latter case, one should also bear in mind that Persides1
$F_{l\omega,+}(r)=F_{l}(\omega r;\epsilon=2i)\sim i^{l+1}\sum_{n=0}^{\infty}\tau_{n}(\omega r)^{-(n+1)}.$ (115)
Appendix E The Poincaré framework
In Ref. Poincare2 , Poincaré studied linear differential equations of the form
$\sum_{k=0}^{n}P_{k}(x)\frac{d^{k}y}{dx^{k}}=0,$ (116)
where the $P_{k}(x)$ are polynomials in the variable $x$. He knew from the work of Fuchs and Thomé that, apart from some exceptional cases, there exist $n$ functions of the following form:
$e^{Q(x)}\;x^{a}\;\left(A_{0}+\frac{A_{1}}{x}+\frac{A_{2}}{x^{2}}+...\right),$ (117)
which satisfy formally Eq. (116), where $Q(x)$ is a polynomial in $x$. The series in Eq. (117) was then said to be a normal series. In particular, if the normal series is convergent, one says that Eq. (116) admits a normal integral.
Poincaré says that a divergent series
$\sum_{k=0}^{\infty}\frac{A_{k}}{x^{k}},$
where the sum of the first $(n+1)$ terms is denoted by $S_{n}$, represents asymptotically a function $J:x\rightarrow J(x)$ if the epression
$x^{n}(J(x)-S_{n}(x))$
approaches 00 when $x$ tends to infinity:
$\lim_{x\to\infty}x^{n}(J(x)-S_{n}(x))=0.$ (118)
This means that, for $x$ suffiently large, one has the majorization
$x^{n}(J(x)-S_{n}(x))<\varepsilon.$ (119)
The error
$J(x)-S_{n}(x)=\frac{\varepsilon}{x^{n}}$ (120)
made on taking (just) the first $(n+1)$ terms is then extremely small, and also much smaller than the error made on taking just $n$ terms, which equals
$J(x)-S_{n-1}(x)=\frac{A_{n}+\varepsilon}{x^{n}},$ (121)
$\varepsilon$ being very small and $A_{n}$ being finite.
The literature on theoretical physics still relies upon the Poincaré definition of asymptotic expansion, whereas in pure mathematics one has also a different, rigorous framework built by Dieudonné Dieudonne .